OpenStudy (anonymous):

Find a ⋅ b. a = 9i + 5j, b = 4i + 3j

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

stack them vertically and multiply them down

OpenStudy (amistre64):

a = 9i + 5j b = 4i + 3j ------------ 36i + 15j

OpenStudy (anonymous):

<36,15>

OpenStudy (amistre64):

correct, which can also be written in many different ways

OpenStudy (anonymous):

Find a ⋅ b. a = <5, 2>, b = <4, 5> would this be <20,10>

OpenStudy (amistre64):

no, 20+10

OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 7>, v = <9, 7>

OpenStudy (anonymous):

what about this?

OpenStudy (amistre64):

youll need the dot product, and the lengths of each vector

OpenStudy (anonymous):

20+10?

OpenStudy (amistre64):

yes, dot product produces a scalar, not a new vector ....

OpenStudy (anonymous):

so 30?

OpenStudy (amistre64):

ignore the i j parts a = 9i + 5j b = 4i + 3j ------------ a.b 36 + 15 = ?

OpenStudy (anonymous):

51

OpenStudy (amistre64):

thats better :)

OpenStudy (anonymous):

okay! so what about Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 7>, v = <9, 7>

OpenStudy (amistre64):

u = <8, 7> ; length? v = <9, 7> ; length? ----------- dot we will need to know these values

OpenStudy (amistre64):

oh, and just as a cursory glance, they are both on the y=7 axis

OpenStudy (anonymous):

-8.3° 1.7° 3.3° 13.3°

OpenStudy (anonymous):

those are the possible answers

OpenStudy (amistre64):

the options are immaterial at the moment u = <8, 7> ; length? v = <9, 7> ; length? ----------- dot im not going to do it for you :)

OpenStudy (anonymous):

i dont even know where to begin

OpenStudy (anonymous):

do i have to multiply agin?

OpenStudy (amistre64):

i showed you how to do that dot, and you should know by this stage how to find the length of a vector

OpenStudy (amistre64):

length is: square, add, sqrt

OpenStudy (anonymous):

so i square the 8 add 7 and then find the sqrt. of that number?

OpenStudy (amistre64):

u = <8, 7> ; sqrt(8^2+7^2) v = <9, 7> ; sqrt(9^2+7^2)

OpenStudy (anonymous):

oh okay

OpenStudy (amistre64):

square the parts, add them together, and sqrt the results gives you length

OpenStudy (anonymous):

the first is 10.63

OpenStudy (amistre64):

now, the angle between them is defined by a formula\[cos(\alpha)=\frac{dot}{length_1*length_2}\]

OpenStudy (anonymous):

second is 11.40

OpenStudy (anonymous):

and the dot is the numbers multiplied right?

OpenStudy (amistre64):

yes, multiply down, and add them up

OpenStudy (anonymous):

121/10.63(11.40)

OpenStudy (amistre64):

lol, add em across might be more visually adept

OpenStudy (anonymous):

is that right?

OpenStudy (amistre64):

72 49 121 good and the lengths are most likely fine as well

OpenStudy (amistre64):

take the arccos, or inverse cosine of that value to determine the angle

OpenStudy (anonymous):

121/121.182

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got .063

OpenStudy (anonymous):

that does not seem right

OpenStudy (amistre64):

i get 0.05778 radians

OpenStudy (amistre64):

or about 3.3 degrees

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

do you have time for two more?

OpenStudy (amistre64):

when you approximate the sqrts, you introduce a small error

OpenStudy (anonymous):

ohh okay!

OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <2, 6>

OpenStudy (amistre64):

dot = 0 defined perp

OpenStudy (anonymous):

do i begin the same way? with the multiplication

OpenStudy (amistre64):

of course, dotting is the same process ....

OpenStudy (anonymous):

okay and if it is 0?

OpenStudy (anonymous):

does that mean it would be neither?

OpenStudy (amistre64):

umm, dot = 0 is perp. orthogonal is another term for perpendicular

OpenStudy (anonymous):

oh okay. I did not know that!!!

OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 8, -6>; v = <3, -8, -3>; w = <-4, -2, -6>

OpenStudy (amistre64):

v.w is still a dot of v and w

OpenStudy (anonymous):

so what is the r?

OpenStudy (amistre64):

extra stuff that isnt used.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so the answer would be 22?

OpenStudy (amistre64):

if i order a burger and fries from McDonalds, the rest of the stuff on the menu is inconsequential

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

v = <3, -8, -3> w = <-4, -2, -6> ------------------ -12+16+18 = ?

OpenStudy (anonymous):

so is 22 right?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

thank you so much for all of your time! i really really really appreciate it!